OpenStudy (anonymous):

What is the current in a parallel circuit which has two resistors (17.2 ohms and 22.4 ohms) and a power source of 6.0 volts?

OpenStudy (anonymous):

6/(17.2+22.4/17.2*22.4)

OpenStudy (mrnood):

 \[\frac{ 1 }{R _{eq} }=\frac{ 1 }{R _{1} }+\frac{ 1 }{ R _{2} }\]